OpenStudy (vane11):

Hi guys, I need help finding A in this equation so that I can go ahead and answer the actual questions (not included) can someone help me? Metro Department Store found that t weeks after the end of a sales promotion the volume of sales was given by S(t) = B + Ae−kt (0 ≤ t ≤ 4) where B = 51,000 and is equal to the average weekly volume of sales before the promotion. The sales volumes at the end of the first and third weeks were $81,790 and $62,240, respectively. Assume that the sales volume is decreasing exponentially.

OpenStudy (campbell_st):

so the equation is \[S(t) = B + Ae^{-kt}\]

OpenStudy (vane11):

yes sorry

OpenStudy (campbell_st):

ok... look at week 1 of the promotion \[81790 = 51000 +Ae^{-k}\] or \[30790 = Ae^{-k}\] and in week 3 \[62240 = 51000 + A e^{-3k}\] or \[10240 = Ae^{-3k}\] so I would split the week 3 equation to \[10240 = Ae^{-k} \times e^{-2k}\] then substitute the week 1 simplified equation. this will allow you to solve for k... then when you have k, you can find A hope it helps

OpenStudy (vane11):

ohhhh ok, thanks!